# RoboMamba: Multimodal State Space Model for Efficient Robot Reasoning and Manipulation

### v1

@@ Abstract
A fundamental objective in robot manipulation is to enable models to comprehend visual scenes and execute actions. Although existing robot Multimodal Large Language Models (MLLMs) can handle a range of basic tasks, they still face challenges in two areas: 1) inadequate reasoning ability to tackle complex tasks, and 2) high computational costs for MLLM fine-tuning and inference. The recently proposed state space model (SSM) known as Mamba demonstrates promising capabilities in non-trivial sequence modeling with linear inference complexity. Inspired by this, we introduce RoboMamba, an end-to-end robotic MLLM that leverages the Mamba model to deliver both robotic reasoning and action capabilities, while maintaining efficient fine-tuning and inference. Specifically, we first integrate the vision encoder with Mamba, aligning visual data with language embedding through co-training, empowering our model with visual common sense and robot-related reasoning. To further equip RoboMamba with action pose prediction abilities, we explore an efficient fine-tuning strategy with a simple policy head. We find that once RoboMamba possesses sufficient reasoning capability, it can acquire manipulation skills with minimal fine-tuning parameters (0.1% of the model) and time (20 minutes). In experiments, RoboMamba demonstrates outstanding reasoning capabilities on general and robotic evaluation benchmarks. Meanwhile, our model showcases impressive pose prediction results in both simulation and real-world experiments, achieving inference speeds 7 times faster than existing robot MLLMs. Our project web page: https://sites.google.com/view/robomamba-web
@@ 1 Introduction
The scaling up of data has significantly propelled research on Large Language Models (LLMs) [1, 2, 3], showcasing notable advancements in reasoning and generalization abilities within Natural Language Processing (NLP). To comprehend multimodal information, Multimodal Large Language Models (MLLMs) [4, 5, 6, 7, 8] have been introduced, empowering LLMs with the capability of visual instruction-following and scene understanding. Inspired by the strong capabilities of MLLMs in general settings, recent research aims to incorporate MLLMs into robot manipulation. On one hand, some works [9, 10, 11, 12] enable robots to comprehend natural language and visual scenes, automatically generating task plans. On the other hand, [13, 14, 15] effectively leverage the inherent capabilities of MLLMs, empowering them with the ability to predict manipulation poses.
@@
Robot manipulation involves interacting with objects in dynamic environments, requiring human-like reasoning abilities to comprehend the semantic information of scenes [11, 16], alongside a robust low-level action prediction ability [17, 18]. While existing MLLM-based approaches can handle a range of basic tasks, they still face challenges in two aspects. First, the reasoning capabilities of pre-trained MLLMs [6, 19] in robotic scenarios are found to be insufficient. As shown in Figure 1 (reasoning example), this deficiency presents challenges for fine-tuned robot MLLMs when they encounter complex reasoning tasks. Second, fine-tuning MLLMs and using them to generate robot manipulation actions incurs higher computational costs due to their expensive attention-based LLMs [20, 21]. To balance the reasoning ability and efficiency, several studies [22, 23, 24] have emerged in the field of NLP. Notably, Mamba [25] introduces the innovative selective State Space Model (SSM), promoting context-aware reasoning while maintaining linear complexity. Drawing inspiration from this, we raise a question: “Can we develop an efficient robot MLLM that possesses strong reasoning capabilities while also acquiring robot manipulation skills in a very cost-effective manner?"
@@
To address this, we propose RoboMamba, an end-to-end robotic MLLM that fully leverages the efficiency of Mamba to achieve robust robotic reasoning and action capabilities. As shown in Figure 1, we initially integrate a vision encoder (e.g., CLIP [26]) with Mamba to empower RoboMamba with visual common sense and robot-related reasoning. Specifically, we proceed with alignment pre-training, activating the cross-modal connector [4, 19] to convert visual information into Mamba’s token embeddings. We then unfreeze Mamba for instructions co-training, utilizing its powerful sequence modeling to comprehend high-level robotic and general instruction data. On top of this, to equip RoboMamba with action pose prediction abilities, we explore an efficient fine-tuning strategy with a simple policy head. Notably, we discover that once RoboMamba possesses sufficient reasoning capabilities, it can acquire pose prediction skills with minimal parameter fine-tuning. The fine-tuned policy head constitutes only 0.1% of the model parameters, which is 10 times smaller than existing robot MLLM [15, 14]. In this way, RoboMamba can simultaneously generate robot reasoning using language responses and predict end-effector poses via the policy head.
@@
To systematically evaluate our end-to-end RoboMamba, we conduct extensive experiments in both simulation and real-world scenarios. First, we assess our reasoning abilities on general and robotic evaluation benchmarks. As shown in Figure 1, RoboMamba, with only 3.2B parameters, achieves competitive performance on several MLLM benchmarks and also delivers promising results on RoboVQA (36.3 BLEU-4) [27]. With its strong reasoning abilities, RoboMamba achieves state-of-the-art (SOTA) manipulation performance in the SAPIEN simulation [28], requiring only a 7MB policy head and less than 20 minutes of fine-tuning on a single A100 GPU. Moreover, RoboMamba achieves an inference speed that is 7 times faster than previous SOTA robot MLLM [15]. Additionally, we evaluate RoboMamba in real-world scenarios, where it can generate long-horizon planning and predict the end-effector pose for each atomic task. In summary, our contributions are as follows:
@@
We innovatively integrate a vision encoder with the efficient Mamba language model to construct our end-to-end RoboMamba, which possesses visual common sense and robot-related reasoning abilities.
@@ 2 Related work
State Space Models (SSMs). SSMs have become effective substitutes for transformers and CNNs due to their linear scalability with sequence length [29, 23]. Recent works [22, 30, 31] use the state space to robustly establish dependencies across long sequences. Especially, Mamba [25] designs the SSM matrices to be functions of the input, creating a learnable selection mechanism that improves adaptability and reasoning capabilities. [32, 33, 34, 35, 36] expand selective SSMs to vision and video tasks. Furthermore, MambaIR [37] focuses on image restoration, and PanMamba [38] addresses pan-sharpening, while DiS [39] integrates SSMs into diffusion models. These findings demonstrate that Mamba exhibits promising performance and efficiency in various visual downstream tasks. With the emergence of SSMs, we make the first attempt to introduce Mamba to address critical challenges in robotics, which demands efficient action capabilities.
@@
Robot Manipulation. Traditional robotic manipulation employs state-based reinforcement learning [40, 41, 42, 43]. In contrast, [44, 11, 45, 46, 47] use state with visual observation as input and then make predictions. Specifically, Where2Act [48] takes visual observations and predicts on actionable pixels and movable regions in objects. Flowbot3d [44] predicts point-wise motion flow on 3D objects. Anygrasp [17] employs point cloud data to learn grasp poses on a large scale datasets. Inspired by the success of MLLMs in general scenarios [49, 50, 19, 51, 52], several studies [13, 16] explore utilizing their common sense reasoning capabilities to address manipulation problems. Palm-E [10] integrates multimodal encodings with LLMs, training them end-to-end for manipulation planning. VoxPoser [11] extracts affordances and constraints from MLLMs to further zero-shot predict trajectories. RoboFlamingo [14] fine-tunes MLLM on vision language manipulation dataset to complete language-conditioned manipulation tasks. ManipLLM [15] introduces specific training scheme for manipulation tasks that equips MLLMs with the ability to predict end-effector poses. ManipVQA [53], enhancing robotic manipulation with physically grounded information processed by MLLM. In this paper, instead of fine-tuning a pre-trained MLLM, we introduce a novel end-to-end MLLM that possesses both robot-related reasoning and pose prediction capabilities. Finally, due to space limitation, we provide additional related work in Appendix F.
@@ 3 RoboMamba
In Section 3.1, we introduce the preliminaries of our proposed robot MLLM, including the problem statement and a description of the language model. Subsequently, in Section 3.2 and 3.3, we describe the architecture of RoboMamba and how we empower it with common sense and robot-related reasoning. In Section 3.4, we outline our proposed robot manipulation fine-tuning strategy, which equips our model with pose prediction skills by minimal fine-tuning parameters and time.
@@ 3.2 RoboMamba architecture
To equip RoboMamba with both visual reasoning and manipulation abilities, we start from pre-trained Large Language Models (LLMs) [25] and visual models to construct an effective MLLM architecture. As shown in Figure 2, we utilize the CLIP visual encoder [26] to extract visual features from input images , where and represent batch size and tokens, respectively. In contrast to [55, 56], we do not adopt the vision encoder ensemble technique, which employs various backbones (i.e., DINOv2 [57], CLIP-ConvNeXt [58], CLIP-ViT) for image feature extraction. The ensemble introduces additional computational costs that severely impact the practicality of robot MLLMs in the real world. Therefore, we demonstrate that a simple and straightforward model design can also achieve strong reasoning abilities when combined with high-quality data and appropriate training strategies. To enable the LLM to understand visual features, we connect the vision encoder to the LLM using a multilayer perceptron (MLP). Through this simple cross-modal connector, RoboMamba can convert visual information into language embedding space . Note that model efficiency is crucial in the field of robotics, as robots need to respond quickly based on human instructions. Therefore, we select Mamba as our language model due to its context-aware reasoning ability and linear computational complexity. Text prompts are encoded into embedding space using the pre-trained tokenizer, then concatenated () with visual tokens and input into Mamba. We leverage Mamba’s powerful sequence modeling to comprehend multimodal information and utilize effective training strategies to develop visual reasoning capabilities (as described in the next section). The output tokens are then detokenized () to produce responses in natural language . The model forward can be represented as follows:
To equip our model with both reasoning and manipulation abilities, we meticulously design a comprehensive training pipeline, which is divided into two stages. We introduce the training recipes of Stage 1 in Section 3.3 and present the robot manipulation fine-tuning in Section 3.4.
3.3 General and robot-related training
@@
After constructing the RoboMamba architecture, the next goal is to train our model to learn general vision and robot-related reasoning abilities. As shown in Figure 2, we divide our Stage 1 training into two steps: alignment pre-training (Stage 1.1) and instruction co-training (Stage 1.2). Specifically, unlike previous MLLM training methods [19, 59, 55], we aim to enable RoboMamba to comprehend both common vision and robotic scenes. Given that the robotics field involves numerous complex and novel tasks, RoboMamba requires enhanced generalization capabilities. Therefore, we adopt a co-training strategy in Stage 1.2, combining high-level robotic data (e.g., task planning) with general instruction data. We find that co-training not only leads to more generalizable robot policies but also enhances general scene reasoning abilities due to the complex reasoning tasks embedded in the robotic data (demonstrated in Appendix C). The training details are shown below:
@@
Stage 1.1: Alignment pre-training. We adopt LLaVA [4] filtered 558k image-text paired dataset for our cross-modal alignment. As shown in Figure 2, we freeze the parameters of the CLIP encoder and Mamba language model, and only update the project layer. In this way, we can align image features with the pre-trained Mamba word embedding.
@@
State 1.2: Instruction co-training. In this stage, we first follow previous MLLM works [4, 55, 56] for general vision instruction data collection. We adopt the 655K LLaVA mixed instruction dataset [4] and 400K LRV-Instruct dataset [60], which aim to learn visual instruction following and mitigate hallucination, respectively. Note that mitigating hallucination plays an important role in robotic scenarios, as the robot MLLM needs to generate task planning based on real scenes instead of imagined ones. For example, existing MLLMs might formulaically answer “open the microwave” with “step 1: find the handle,” but many microwaves do not have handles. Next, we incorporate the 800K RoboVQA dataset [27] to learn high-level robotic skills, such as long-horizon planning, success classification, discriminative and generative affordance, past description, and future prediction. During co-training, as shown in Figure 2, we freeze the parameters of the CLIP encoder and fine-tune the projection layer and Mamba on the combined 1.8 million datasets. All outputs from the Mamba language model are supervised using the cross-entropy loss.
@@ 3.4 Robot manipulation fine-tuning
Building upon RoboMamba’s strong reasoning ability, we introduce our robot manipulation fine-tuning strategy in this section, termed Training Stage 2 in Figure 2. Existing MLLM-based manipulation methods [15, 14] require updating the projection layer and the entire LLM during the manipulation fine-tuning stage. While this paradigm can develop action prediction capabilities, it also breaks the inherent abilities of the MLLM and demands significant training resources. To address these challenges, we propose an efficient fine-tuning strategy, as shown in Figure 2. We freeze all the parameters of RoboMamba and introduce a simple policy head to model Mamba’s output tokens. The policy head contains two MLPs that separately learn the end-effector’s position and direction , collectively occupying 0.1% of the entire model parameters. Following [48], the position and direction losses are formulated as follows:
where represents the number of training samples, means the trace of matrix . RoboMamba only predicts the 2D position (, ) of the contact pixel in the image, which is then translated into 3D space using depth information. To evaluate this fine-tuning strategy, we generate a dataset of 10k end-effector pose predictions using the SAPIEN simulation [28]. After manipulation fine-tuning, we find that once RoboMamba possesses sufficient reasoning capabilities, it can acquire pose prediction skills with extremely efficient fine-tuning. Due to the minimal fine-tuning parameters (7MB) and efficient model design, we need only 20 minutes to achieve novel manipulation skills. This finding highlights the importance of reasoning abilities for learning manipulation skills and presents a new perspective: we can efficiently equip an MLLM with manipulation abilities without compromising its inherent reasoning capabilities. Finally, RoboMamba can use language responses for common sense and robot-related reasoning, and the policy head for action pose prediction.

### v2

# RoboMamba: Efficient Vision-Language-Action Model for Robotic Reasoning and Manipulation

@@ Abstract
A fundamental objective in robot manipulation is to enable models to comprehend visual scenes and execute actions. Although existing Vision-Language-Action (VLA) models for robots can handle a range of basic tasks, they still face challenges in two areas: (1) insufficient reasoning ability to tackle complex tasks, and (2) high computational costs for VLA model fine-tuning and inference. The recently proposed state space model (SSM) known as Mamba demonstrates promising capabilities in non-trivial sequence modeling with linear inference complexity. Inspired by this, we introduce RoboMamba, an end-to-end robotic VLA model that leverages Mamba to deliver both robotic reasoning and action capabilities, while maintaining efficient fine-tuning and inference. Specifically, we first integrate the vision encoder with Mamba, aligning visual tokens with language embedding through co-training, empowering our model with visual common sense and robotic-related reasoning. To further equip RoboMamba with SE(3) pose prediction abilities, we explore an efficient fine-tuning strategy with a simple policy head. We find that once RoboMamba possesses sufficient reasoning capability, it can acquire manipulation skills with minimal fine-tuning parameters (0.1% of the model) and time. In experiments, RoboMamba demonstrates outstanding reasoning capabilities on general and robotic evaluation benchmarks. Meanwhile, our model showcases impressive pose prediction results in both simulation and real-world experiments, achieving inference speeds 3 times faster than existing VLA models. Our project web page: https://sites.google.com/view/robomamba-web
@@ 1 Introduction
The scaling up of data has significantly propelled research on Large Language Models (LLMs) [1, 2, 3], showcasing notable advancements in reasoning and generalization abilities within Natural Language Processing (NLP). To comprehend multimodal information, Multimodal Large Language Models (MLLMs) [4, 5, 6, 7, 8] have been introduced, empowering LLMs with the capability of visual instruction-following and scene understanding. Inspired by the strong capabilities of MLLMs in general settings, recent research aims to incorporate MLLMs into robot manipulation. On the one hand, some works [9, 10, 11, 12] enable robots to comprehend natural language and visual scenes, automatically generating task plans. On the other hand, Vision-Language-Action (VLA) models [13, 14, 15] leverage the inherent capabilities of MLLMs, empowering them with the ability to predict low-level SE(3) poses.
@@
Robot manipulation involves interacting with objects in dynamic environments, requiring human-like reasoning abilities to comprehend the semantic information of scenes [11, 16], alongside a robust low-level action prediction ability [17, 18]. While existing MLLM-based policies can handle a range of basic tasks, they still face challenges in two aspects. First, the reasoning capabilities of pre-trained MLLMs [6, 19] in robotic scenarios are found to be insufficient. As shown in Figure 1 (reasoning example), this deficiency presents challenges for fine-tuned robot MLLMs when they encounter complex reasoning tasks. Second, fine-tuning MLLMs and using them to generate robot manipulation actions incurs higher computational costs due to their expensive attention-based LLMs [20, 21]. To balance the reasoning ability and efficiency, several studies [22, 23, 24] have emerged in the field of NLP. Notably, Mamba [25] introduces the innovative selective State Space Model (SSM), promoting context-aware reasoning while maintaining linear complexity. Drawing inspiration from this, we raise a question: “Can we develop an efficient robotic VLA model that possesses strong reasoning capabilities while also acquiring robot manipulation skills in a cost-effective manner?"
@@
To address this, we propose RoboMamba, an efficient robotic VLA that empowers the Mamba LLM to achieve robust robotic reasoning and action capabilities. As shown in Figure 1, we initially integrate a vision encoder (e.g., CLIP [26]) with Mamba to empower RoboMamba with visual common sense and robotic-related reasoning. Specifically, we proceed with alignment pre-training, activating the cross-modal connector [4, 19] to convert visual information into Mamba’s token embeddings. We then unfreeze Mamba for instructions co-training, utilizing its powerful sequence modeling to comprehend high-level robotic and general instruction data. On top of this, to equip RoboMamba with SE(3) pose prediction abilities, we explore an efficient fine-tuning strategy with a simple policy head. Notably, we discover that once RoboMamba possesses sufficient reasoning capabilities, it can acquire pose prediction skills with minimal parameter fine-tuning. The fine-tuned policy head constitutes only 0.1% of the model parameters, which is 10 times smaller than existing robotic VLA approaches [15, 14]. In this way, RoboMamba can simultaneously generate robot reasoning using language responses and predict end-effector poses via the policy head.
@@
To systematically evaluate our proposed RoboMamba, we conduct extensive experiments in both simulation and real-world scenarios. First, we assess our reasoning abilities on general and robotic evaluation benchmarks. RoboMamba, with only 3.2B parameters, achieves competitive performance on several MLLM benchmarks and also delivers promising results on RoboVQA (42.8 BLEU-4) [27]. With its strong reasoning abilities, RoboMamba achieves state-of-the-art (SOTA) manipulation performance in the SAPIEN simulation [28], requiring only a 7MB policy head and a few dozen minutes of fine-tuning on a single A100 GPU. Moreover, RoboMamba achieves an inference speed that is 3 times faster than previous robotic VLA models [29, 15]. Additionally, we evaluate RoboMamba in real-world scenarios, where it can generate long-horizon planning and predict the end-effector pose for each atomic task. In summary, our contributions are as follows:
@@
We introduce RoboMamba, an efficient VLA model that integrates a vision encoder with the linear-complexity Mamba LLM, which possesses visual common sense and robotic-related reasoning abilities.
@@ 2 Related work
State Space Models (SSMs). SSMs have become effective substitutes for transformers and CNNs due to their linear scalability with sequence length [30, 23]. Recent works [22, 31, 32] use the state space to robustly establish dependencies across long sequences. Especially, Mamba [25] designs the SSM matrices to be functions of the input, creating a learnable selection mechanism that improves adaptability and reasoning capabilities. [33, 34, 35, 36, 37, 38] expand selective SSMs to vision and video tasks. Furthermore, MambaIR [39] focuses on image restoration, and PanMamba [40] addresses pan-sharpening, while DiS [41] integrates SSMs into diffusion models. These findings demonstrate that Mamba exhibits promising performance and efficiency in various visual downstream tasks. With the emergence of SSMs, we make the first attempt to introduce Mamba to address critical challenges in robotics, which demands efficient action capabilities.
Multimodal Large Language Models. Large language models (LLMs) have exhibited remarkable reasoning capabilities across various downstream tasks [19, 42, 2]. When addressing complex multimodal reasoning challenges, multimodal large language models (MLLMs) have shown exceptional visual understanding, i.e., BLIP-2 [43], OpenFlamingo [44], LLaMA-Adapter [19, 45], and LLaVA [46]. Additionally, the introduction of 3D MLLMs [12, 47, 48] seeks to expand the reasoning and conversational capabilities of LLMs to include the 3D modality. However, deploying LMMs is expensive due to their significant computational overhead, primarily caused by their billions of parameters. To mitigate these challenges, recent small-scale models [49, 50] demonstrate impressive performance while maintaining manageable computational costs. LLaVA-Phi [49] empowers the recently developed smaller LLM, Phi-2, for visual instruction tuning. TinyLLaVA [50] and MobileVLM V2 [51] demonstrate that high-quality training data and schemes can effectively compensate for the reasoning abilities of smaller LMMs. Furthermore, Cobra [52] innovatively utilizes an SSM-based Mamba LLM to reduce complexity and improve inference speed on common sense reasoning tasks. Different from previous works, our goal is to develop an efficient Robotic VLA model using the SSM-based language model. This model not only possesses common sense understanding but also has the capability to complete manipulation tasks effectively.
@@
Robot Manipulation. Traditional robotic manipulation employs state-based reinforcement learning [53, 54, 55, 56]. In contrast, [57, 11, 58, 59, 60] use state with visual observation as input and then make predictions. Specifically, Where2Act [61] takes visual observations and predicts on actionable pixels and movable regions in objects. Flowbot3d [57] predicts point-wise motion flow on 3D objects. Anygrasp [17] employs point cloud data to learn grasp poses on a large scale datasets. Inspired by the success of MLLMs in general scenarios [43, 44, 19, 45, 46], several VLA models [13, 16] explore utilizing their common sense reasoning capabilities to address manipulation problems. Palm-E [10] integrates multimodal encodings with LLMs, training them end-to-end for manipulation planning. VoxPoser [11] extracts affordances and constraints from MLLMs to further zero-shot predict trajectories. RoboFlamingo [14] fine-tunes MLLM on vision language manipulation dataset to complete language-conditioned manipulation tasks. ManipLLM [15] introduces specific training scheme for manipulation tasks that equips MLLMs with the ability to predict end-effector poses. ManipVQA [62], enhancing robotic manipulation with physically grounded information processed by MLLM. In this paper, instead of fine-tuning a pre-trained MLLM, we introduce a novel efficient VLA model that possesses both robotic-related reasoning and low-level pose prediction capabilities.
@@ 3 RoboMamba
In Section 3.1, we introduce the preliminaries of our proposed RoboMamba, including the problem statement and a description of the language model. Subsequently, in Section 3.2 and 3.3, we describe the architecture of RoboMamba and how we empower it with common sense and robotic-related reasoning. In Section 3.4, we outline our proposed robot manipulation fine-tuning strategy, which equips our model with pose prediction skills by minimal fine-tuning parameters and time.
@@ 3.2 RoboMamba architecture
To equip RoboMamba with both visual reasoning and manipulation abilities, we start from pre-trained Large Language Models (LLMs) [25] and visual models to construct an effective VLA model architecture. As shown in Figure 2, we utilize the CLIP visual encoder [26] to extract visual features from input images , where and represent batch size and tokens, respectively. In contrast to [64, 52], we do not adopt the vision encoder ensemble technique, which employs various backbones (i.e., DINOv2 [65], CLIP-ConvNeXt [66], CLIP-ViT) for image feature extraction. The ensemble introduces additional computational costs that severely impact the practicality of VLA model in the real world. Therefore, we demonstrate that a simple and straightforward model design can also achieve strong reasoning abilities when combined with high-quality data and appropriate training strategies. To enable the LLM to understand visual features, we connect the vision encoder to the LLM using a multilayer perceptron (MLP). Through this simple cross-modal connector, RoboMamba can convert visual information into language embedding space . Note that model efficiency is crucial in the field of robotics, as robots need to respond quickly based on human instructions. Therefore, we select Mamba as our language model due to its context-aware reasoning ability and linear computational complexity. Text prompts are encoded into embedding space using the pre-trained tokenizer, then concatenated () with visual tokens and input into Mamba. We leverage Mamba’s powerful sequence modeling to comprehend multimodal information and utilize effective training strategies to develop visual reasoning capabilities (as described in the next section). The output tokens are then detokenized () to produce responses in natural language . To equip our model with both reasoning and manipulation abilities, we meticulously design a comprehensive training pipeline, which is divided into two stages. We introduce the training recipes of Stage 1 in Section 3.3 and present the robot manipulation fine-tuning in Section 3.4.
3.3 General and robotic-related training
@@
After constructing the RoboMamba architecture, the next goal is to train our model to learn general vision and robotic-related reasoning abilities. As shown in Figure 2, we divide our Stage 1 training into two steps: alignment pre-training (Stage 1.1) and instruction co-training (Stage 1.2). Specifically, unlike previous MLLM training methods [19, 67, 64], we aim to enable RoboMamba to comprehend both common vision and robotic scenes. Given that the robotics field involves numerous complex and novel tasks, RoboMamba requires enhanced generalization capabilities. Therefore, we adopt a co-training strategy in Stage 1.2, combining high-level robotic data (e.g., task planning) with general instruction data. We find that co-training not only leads to more generalizable robotic policies but also enhances the majority of general scene reasoning abilities due to the complex reasoning tasks embedded in the robotic data (demonstrated in Appendix C). The training details are shown below:
@@
Stage 1.1: Alignment pre-training. We adopt LLaVA [4] filtered 558k image-text paired dataset for our cross-modal alignment. As shown in Figure 2, we freeze the parameters of the vision encoder and Mamba language model, and only update the project layer. In this way, we can align image features with the pre-trained Mamba word embedding.
@@
State 1.2: Instruction co-training. In this stage, we first follow previous MLLM works [4, 64, 52] for general vision instruction data collection. We adopt the 655K LLaVA mixed instruction dataset [4], the ShareGPT4V-SFT dataset [68], or the LLaVA-Next dataset [69]. In our dataset selection, mitigating hallucination is crucial in robotic scenarios, as the robotic MLLM needs to generate task planning based on real scenes rather than imagined ones. For example, existing MLLMs might formulaically answer “open the microwave” with “step 1: find the handle,” but many microwaves do not have handles. Next, we incorporate the RoboVQA dataset [27] to learn high-level robotic skills, such as long-horizon planning, success classification, discriminative and generative affordance, past description, and future prediction. During co-training, as shown in Figure 2, we freeze the parameters of the CLIP encoder and fine-tune the projection layer and Mamba on the combined instruction datasets. All outputs from the Mamba language model are supervised using the cross-entropy loss.
@@ 3.4 Robot manipulation fine-tuning
Building upon RoboMamba’s strong reasoning ability, we introduce our robot manipulation fine-tuning strategy in this section, termed Training Stage 2 in Figure 2. Existing manipulation VLA models [29, 15, 14] require updating the projection layer and the LLM during the manipulation fine-tuning stage. While this paradigm can develop action prediction capabilities, it also breaks the inherent abilities of the pre-trained model and demands significant training resources. To address these challenges, we propose an efficient fine-tuning strategy, as shown in Figure 2. We freeze all the parameters of RoboMamba and introduce a simple policy head to model Mamba’s output tokens. The policy head contains two types of MLPs that separately learn the end-effector’s position and direction , collectively occupying around 0.1% of the model’s total parameters. Following [61], the position and direction losses are formulated as follows:
where represents the number of training samples, means the trace of matrix . In our open-loop simulation experiments, RoboMamba only predicts the 2D position (, ) of the contact pixel in the image, which is then translated into 3D space using depth information. We also derive the gripper’s left direction (gripper z-forward) from its up and forward orientations based on geometric relationships. To evaluate this fine-tuning strategy, we generate a dataset of 10k end-effector pose predictions using the SAPIEN simulation [28]. After manipulation fine-tuning, we find that once RoboMamba possesses sufficient reasoning capabilities, it can acquire pose prediction skills with extremely efficient fine-tuning. Due to the minimal fine-tuning parameters (7MB) and efficient model design, we need only a few dozen minutes to achieve novel manipulation skills. This finding highlights the importance of reasoning abilities for learning manipulation skills and presents a new perspective: we can efficiently equip an VLA model with manipulation abilities without compromising its inherent reasoning capabilities. Finally, RoboMamba can use language responses for common sense and robotic-related reasoning, and the policy head for action pose prediction.